OpenStudy (anonymous):

simplify b^2(a+2)/b(a+2)^2

OpenStudy (anonymous):

b/a+2

OpenStudy (saifoo.khan):

\[\frac{b^2(a+2)}{b(a+2)^2}\] \[\frac{b*b(a+2)}{b(a+2)(a+2)}\] Cancel out "b" \[\frac{b(a+2)}{(a+2)(a+2)}\] Now cancel out (a+2) from both sides \[\frac{b}{(a+2)}\]

OpenStudy (anonymous):

Thanks.

OpenStudy (saifoo.khan):

Welcome.